IR-MICROWAVE DOUBLE RESONANCE STUDIES OF DIPOLE MOMENTS IN THE ν_{1} and ν_{3} STATES OF AMMONIA

\underline{MARK D. MARSHALL}, Department of Chemistry-2243, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000; KAMIL CAN IZGI and J. S. MUENTER, Department of Chemistry, University of Rochester, Rochester, NY 14627.

Infrared laser-microwave double resonance spectroscopy is used to observe the Stark effect of tunneling transitions within the ν_{1} and ν_{3} hydrogen stretching states of NH_{3}. Dipole moments for 11 J,K states of ν_{1} and 3 J,K states of ν_{3} are measured to high accuracy. These data, combined with previous measurements in the ground and excited bending states, give a dipole moment function of:

μ_{v1v2v3v4} = 1.5610 + 7.2 × 10^{−3}(v_{1} + \frac{1}{2} ) − 2.271 × 10^{−1}(v_{2} + \frac{1}{2} ) + 3.75 × 10^{−2}(v_{3} + 1) − 1.65 × 10^{−2}(v_{4} + 1).

μ_{e} = 1.561 D is the first experimental measurement of the NH_{3} equilibrium moment. These results are also used to analyze and recalculate the dipole moment measured by Shimizu and coworkers ^{a} in a v_{NH} = 5 excited state of NH_{3}.

aK. Nakawa, Y. Moriwaki, and T. Shimizu, Opt. Lett. 14, 488 (1989).